# Approximate independence of permutation mixtures

### v2

@@ Appendix C Deferred proofs
Here $\delta_{x}$ is the Dirac delta measure on the singleton $\{x\}$. It is clear that ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})\leq|{\mathcal{P}}|-1<{\mathsf{C}}$ (cf. Lemma C.1), and $\mathsf{\Delta}_{H^{2}}({\mathcal{P}})=\frac{1}{\Delta}$. Now consider $\chi^{2}(\mathbb{P}_{mn}\|\mathbb{Q}_{mn})$, with each $P_{i}$ appearing $n$ times. By Lemma 4.1,
@@
C.12 Proof of Lemma 5.3
Let $\delta=\delta(\varepsilon)>0$ be any fixed constant such that
$(1+\delta)(1-\varepsilon)<1-\frac{\varepsilon}{2}.$ (46)
$I(\theta;X^{\lambda})=\frac{1}{2}\int_{0}^{\lambda}\mathbb{E}\left[\|\theta-\mathbb{E}\left[\theta|X^{t}\right]\!\|^{2}\right]\mathrm{d}t=:\frac{1}{2}\int_{0}^{\lambda}\mathrm{mmse}(t)\mathrm{d}t.$
We denote by $I_{\mathbb{P}}(\theta;X^{\lambda})$ and $I_{\mathbb{Q}}(\theta;X^{\lambda})$ the mutual information under the permutation prior and the i.i.d. prior, respectively. Similarly we use the notations $(\mathrm{mmse}_{\mathbb{P}}(t),\mathrm{mmse}_{\mathbb{Q}}(t))$ for the minimum mean squared errors (MMSEs) under these priors. By the I-MMSE formula,
For the difference in the mutual information, the inequality (1) in Corollary 1.4 states that
$0\leq I_{\mathbb{Q}}(\theta;X^{\lambda})-I_{\mathbb{P}}(\theta;X^{\lambda})\leq 10\sum_{\ell=2}^{n}I_{\chi^{2}}(\theta_{1};X_{1}^{\lambda})^{\ell}.$
Consequently, using the decreasing property of $t\mapsto\mathrm{mmse}(t)$ under both priors, we have
which rearranges to
In the above inequality, note that $\mathrm{mmse}_{\mathbb{P}}(1)$ is precisely the target Bayes risk under the permutation prior. As for $\mathrm{mmse}_{\mathbb{Q}}(1+\delta)$, this quantity corresponds to the Bayes risk under a rescaled i.i.d. prior $\left((1-\alpha)\delta_{0}+\alpha\delta_{\mu_{0}}\right)^{\otimes n}$ for the Gaussian mean vector, with
$\mu_{0}:=\sqrt{(1+\delta)(1-\varepsilon)}\mu=\sqrt{2(1+\delta)(1-\varepsilon)\log(1/R^{p})},\qquad\alpha:=\frac{\lfloor s\rfloor}{n}.$
As $\mu_{0}<\mu$ by (46) and $s=\omega(1)$, this Bayes risk has been characterized in [DJ94, Section 6] that $\mathrm{mmse}_{\mathbb{Q}}(1+\delta)=(1-o(1))s\mu_{0}^{2}>(1-o(1))(1-\varepsilon)s\mu^{2}$. Finally, since $\delta=\delta(\varepsilon)>0$ is a fixed constant and $s\mu^{2}=\Omega(1)$, the inequality (47) proves the corollary as long as we show that
$I_{\chi^{2}}(\theta_{1};X_{1}^{1+\delta})=o_{\varepsilon}(1).$ (48)
[…]
Consequently,
$I_{\chi^{2}}(\theta_{1};X_{1}^{1+\delta})$ =$\overset{\rm(a)}{=}\alpha(1-\alpha)\int\frac{(\mathrm{d}{\mathcal{N}}(\mu_{0},1)-\mathrm{d}{\mathcal{N}}(0,1))^{2}}{\mathrm{d}[\alpha{\mathcal{N}}(\mu_{0},1)+(1-\alpha){\mathcal{N}}(0,1)]}$
=$\overset{\rm(b)}{=}\alpha(1-\alpha)\mathbb{E}_{Z\sim{\mathcal{N}}(0,1)}\left[\frac{\left(\exp(\mu_{0}Z-\mu_{0}^{2}/2)-1\right)^{2}}{\alpha\exp(\mu_{0}Z-\mu_{0}^{2}/2)+1-\alpha}\right],$
$\mathbb{E}_{Z\sim{\mathcal{N}}(0,1)}\left[\frac{\left(\exp(\mu_{0}Z-\mu_{0}^{2}/2)-1\right)^{2}}{\alpha\exp(\mu_{0}Z-\mu_{0}^{2}/2)+1-\alpha}{\mathbf{1}_{\left\{{Z\leq(1+\frac{\varepsilon}{4})\mu_{0}}\right\}}}\right]$
$\leq\frac{1}{1-\alpha}\left(1+\mathbb{E}_{Z\sim{\mathcal{N}}(0,1)}\left[\exp(2\mu_{0}Z-\mu_{0}^{2}){\mathbf{1}_{\left\{{Z\leq(1+\frac{\varepsilon}{4})\mu_{0}}\right\}}}\right]\right)$
=$\overset{\rm(c)}{=}\frac{1}{1-\alpha}\left(1+e^{\mu_{0}^{2}}\cdot\mathbb{P}\left({\mathcal{N}}(2\mu_{0},1)\leq\left(1+\frac{\varepsilon}{4}\right)\mu_{0}\right)\right)$
$\overset{\rm(d)}{\leq}\frac{1}{1-\alpha}\left(1+\exp\left(\mu_{0}^{2}-\frac{1}{2}\left(1-\frac{\varepsilon}{4}\right)^{2}\mu_{0}^{2}\right)\right)=\frac{1}{1-\alpha}\left(1+\exp\left(\left(\frac{1}{2}+\frac{\varepsilon}{4}-\frac{\varepsilon^{2}}{32}\right)\mu_{0}^{2}\right)\right),$
$\mathbb{E}_{Z\sim{\mathcal{N}}(0,1)}\left[\frac{\left(\exp(\mu_{0}Z-\mu_{0}^{2}/2)-1\right)^{2}}{\alpha\exp(\mu_{0}Z-\mu_{0}^{2}/2)+1-\alpha}{\mathbf{1}_{\left\{{Z>(1+\frac{\varepsilon}{4})\mu_{0}}\right\}}}\right]$
$\leq\frac{1}{\alpha}\mathbb{E}_{Z\sim{\mathcal{N}}(0,1)}\left[\exp(\mu_{0}Z-\mu_{0}^{2}/2){\mathbf{1}_{\left\{{Z>(1+\frac{\varepsilon}{4})\mu_{0}}\right\}}}\right]$
$=\frac{1}{\alpha}\mathbb{P}\left({\mathcal{N}}(\mu_{0},1)>\left(1+\frac{\varepsilon}{4}\right)\mu_{0}\right)\leq\frac{1}{\alpha}\exp\left(-\frac{\varepsilon^{2}\mu_{0}^{2}}{32}\right).$
Since $\alpha\leq\exp\left(-\frac{(1+\varepsilon/2)\mu_{0}^{2}}{2}\right)$, we have
$I_{\chi^{2}}(\theta_{1};X_{1}^{1+\delta})\leq\alpha+2\exp\left(-\frac{\varepsilon^{2}\mu_{0}^{2}}{32}\right)=o_{\varepsilon}(1),$
as claimed. The proof of (48) is then complete.
C.13 Proof of Lemma 5.4
$P_{\pi}$ $:=P_{\pi(1)}\otimes\cdots\otimes P_{\pi(n)}$
@@
$\chi^{2}(\overline{\mathbb{P}}\|\mathbb{P}_{-0})$ $=\chi^{2}\left(\mathbb{E}_{\pi\sim\text{\rm Unif}(S_{n})}\left[\frac{1}{n+1}\sum_{j=0}^{n}P_{\pi,j}\right]\bigg{\|}\mathbb{E}_{\pi\sim\text{\rm Unif}(S_{n})}\left[P_{\pi}\right]\right)$
$\leq\mathbb{E}_{\pi\sim\text{\rm Unif}(S_{n})}\left[\chi^{2}\left(\frac{1}{n+1}\sum_{j=0}^{n}P_{\pi,j}\bigg{\|}P_{\pi}\right)\right]$
$\overset{\rm(a)}{\leq}\frac{n}{n+1}\mathbb{E}_{\pi\sim\text{\rm Unif}(S_{n})}\left[\chi^{2}\left(\frac{1}{n}\sum_{j=1}^{n}P_{\pi,j}\bigg{\|}P_{\pi}\right)\right].$
Consequently, for $j\neq j^{\prime}$,
and
Therefore, the second moment method gives
$\chi^{2}\left(\frac{1}{n}\sum_{j=1}^{n}P_{\pi,j}\bigg{\|}P_{\pi}\right)=\frac{1}{n^{2}}\sum_{j,j^{\prime}=1}^{n}\int\frac{\mathrm{d}P_{\pi,j}\mathrm{d}P_{\pi,j^{\prime}}}{\mathrm{d}P_{\pi}}-1\leq\frac{{\mathsf{D}}_{\chi^{2}}({\mathcal{P}})}{n},$
and we obtain the first claim. The second claim follows from
where (b) is the triangle inequality of the Hellinger distance, and (c) uses $H^{2}(P,Q)\leq\chi^{2}(P\|Q)$.
C.14 Proof of Lemma 5.5
Since
@@
$\mathbb{E}\left[\vartheta_{1}|X_{1}\right]-\mathbb{E}\left[\vartheta_{1}|X\right]=\sum_{i=1}^{n}\theta_{i}\left(P_{\vartheta_{1}=\theta_{i}|X_{1}}-P_{\vartheta_{1}=\theta_{i}|X}\right)=\sum_{i=1}^{n}\theta_{i}P_{\vartheta_{1}=\theta_{i}|X_{1}}\left(1-\frac{P_{X_{2}^{n}|\vartheta_{1}=\theta_{i}}}{P_{X_{2}^{n}|X_{1}}}\right)$
according to the Bayes rule, we have
$\overline{P}:=P_{X_{2}^{n}|X_{1}}=\sum_{i=1}^{n}w_{i}P_{i}.$
and therefore
To upper bound the largest eigenvalue of $\overline{A}$, note that Lemma 5.4 implies that $H^{2}(P_{i},P_{j})\leq\frac{4{\mathsf{D}}_{\chi^{2}}({\mathcal{P}})}{n}$ for all $i,j\in[n]$. Consequently, following the spectral gap argument in Lemma 4.2, it holds that
In addition,
$=\frac{1}{n^{2}}\sum_{i=1}^{n}\left(1+\chi^{2}\bigg{(}P_{\theta_{i}}\bigg{\|}\frac{1}{n}\sum_{j=1}^{n}P_{\theta_{j}}\bigg{)}\right)\leq\frac{1+{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})}{n}.$
A combination of the above inequalities shows that
as claimed.

### v3

@@ Appendix C Deferred proofs
Here $\delta_{x}$ is the Dirac delta measure on the singleton $\{x\}$. It is clear that ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})\leq|{\mathcal{P}}|-1<{\mathsf{C}}$ (cf. Lemma˜C.1), and $\mathsf{\Delta}_{H^{2}}({\mathcal{P}})=\frac{1}{\Delta}$. Now consider $\chi^{2}(\mathbb{P}_{mn}\|\mathbb{Q}_{mn})$, with each $P_{i}$ appearing $n$ times. By Lemma˜5.1,
@@
C.17 Proof of Lemma˜6.3
For the capacity upper bound, let $Q_{1},\dots,Q_{m}$ be an $\varepsilon$-cover of ${\mathcal{P}}_{n}$ under the KL divergence, with a fixed $\varepsilon>0$ and $m=N_{\mathrm{KL}}({\mathcal{P}}_{n},\varepsilon)$. By the standard entropic upper bound of the mutual information (cf. [YB99] or [PW25, Theorem 32.4]), one has
$I(X^{n};Y^{n})$ $\leq\mathbb{E}_{X^{n}}\left[\min_{i\in[m]}{\sf KL}(P_{Y^{n}|X^{n}}\|Q_{i}^{\otimes n})\right]+\log m$
@@
$\overset{\rm(b)}{=}\mathbb{E}_{X^{n}}\Bigg{[}{\sf KL}(P_{Y^{n}|X^{n}}\|\overline{P}^{\otimes n})+\min_{i\in[m]}\mathbb{E}_{P_{Y^{n}|X^{n}}}\left[\log\frac{\mathrm{d}\overline{P}^{\otimes n}}{\mathrm{d}Q_{i}^{\otimes n}}\right]\Bigg{]}+\log m$
@@
$\overset{\rm(c)}{=}\mathbb{E}_{\theta}\left[{\sf KL}(P_{Y^{n}|X^{n}}\|\overline{P}^{\otimes n})+\min_{i\in[m]}n\cdot{\sf KL}(\overline{P}\|Q_{i})\right]+\log m$
where $\overline{P}$ denotes the average distribution $\frac{1}{n}\sum_{i=1}^{n}{\mathsf{K}}_{X_{i}}$ in (b), (c) uses $\mathrm{Law}(Y_{i}|X^{n})=\overline{P}$ for all $i\in[n]$, and (d) follows from Theorem˜1.2 and the definition of the KL covering applied to $\overline{P}\in{\mathcal{P}}_{n}$. Taking the infimum over $\varepsilon>0$ then completes the proof of the upper bound.
For the capacity lower bound, let $Q_{1},\dots,Q_{m}$ be an $\varepsilon$-packing of ${\mathcal{P}}_{n}$ under the Hellinger distance, with a fixed $\varepsilon>0$ and $m=M_{\mathrm{H}}({\mathcal{P}}_{n},\varepsilon)$. By the definition of ${\mathcal{P}}_{n}$, each $Q_{i}$ can be expressed as
$Q_{i}=\frac{1}{n}\sum_{j=1}^{n}{\mathsf{K}}_{x_{ij}},\quad\text{for some }x_{i}^{n}:=(x_{i1},\dots,x_{in})\in{\mathcal{X}}^{n}.$
Now let $X^{n}\sim\text{\rm Unif}(\left\{x_{1}^{n},\dots,x_{m}^{n}\right\})$. By [HO97, Lemma 3], it holds that
$I(X^{n};Y^{n})\geq-\frac{1}{m}\sum_{i=1}^{m}\log\Bigg{(}\frac{1}{m}\sum_{j=1}^{m}\exp\left(-\frac{1}{2}D_{1/2}(P_{Y^{n}|X^{n}=x_{i}^{n}},P_{Y^{n}|X^{n}=x_{j}^{n}})\right)\Bigg{)},$ (54)
where $D_{1/2}(P,Q)=-2\log(1-H^{2}(P,Q)/2)=-2\log\int\sqrt{\mathrm{d}P\mathrm{d}Q}$ is the $\frac{1}{2}$-Rényi divergence. To proceed, we prove the following perturbation bound: for any $P^{\prime},Q^{\prime}$, it holds that
$D_{1/2}(P,Q)\geq\frac{1}{2}\left(D_{1/2}(P^{\prime},Q^{\prime})-\log(\chi^{2}(P\|P^{\prime})+1)-\log(\chi^{2}(Q\|Q^{\prime})+1)\right).$ (55)
Indeed, by Hölder’s inequality, we have
$\left(\int\sqrt{\mathrm{d}P\mathrm{d}Q}\right)^{4}\leq\left(\int\sqrt{\mathrm{d}P^{\prime}\mathrm{d}Q^{\prime}}\right)^{2}\left(\int\frac{(\mathrm{d}P)^{2}}{\mathrm{d}P^{\prime}}\right)\left(\int\frac{(\mathrm{d}Q)^{2}}{\mathrm{d}Q^{\prime}}\right),$
so that taking the logarithm gives (55). Next, for $i\neq j$, choosing $P=P_{Y^{n}|X^{n}=x_{i}^{n}},Q=P_{Y^{n}|X^{n}=x_{j}^{n}},P^{\prime}=Q_{i}^{\otimes n}$, and $Q^{\prime}=Q_{j}^{\otimes n}$ in (55) gives
$D_{1/2}(P_{Y^{n}|X^{n}=x_{i}^{n}},P_{Y^{n}|X^{n}=x_{j}^{n}})$ $\geq\frac{1}{2}\left(nD_{1/2}(Q_{i},Q_{j})-2{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})(1+\mathsf{\Delta}_{H^{2}}({\mathcal{P}}))\right)$
$\geq\frac{n\varepsilon^{2}}{2}-{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})(1+\mathsf{\Delta}_{H^{2}}({\mathcal{P}})),$
for $D_{1/2}(Q_{i},Q_{j})\geq H^{2}(Q_{i},Q_{j})\geq\varepsilon^{2}$. Therefore, (54) gives
$I(X^{n};Y^{n})$ $\geq-\log\left(\exp\left(-\frac{n\varepsilon^{2}}{4}+\frac{1}{2}{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})(1+\mathsf{\Delta}_{H^{2}}({\mathcal{P}}))\right)+\frac{1}{m}\right)$
$\geq\min\left\{\log m,\frac{n\varepsilon^{2}}{4}-\frac{1}{2}{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})(1+\mathsf{\Delta}_{H^{2}}({\mathcal{P}}))\right\}-\log 2,$
using $\frac{1}{a}+\frac{1}{b}\leq\frac{2}{\min\{a,b\}}$ for $a,b>0$. Taking the supremum over $\varepsilon>0$ proves the lower bound.